A solution with a pH of 5.0 ______.


A. has a hydrogen-ion concentration of 5.0 M


B. has a hydrogen-ion concentration of 1 x 10-5 M


C. is basic


D. is neutral

1
Expert's answer
2022-04-11T17:05:03-0400

pH = -log[H+];

[H+] = 10^(-pH) = 10^(-5) M.

This solution is acid.

So, correct answer is B (has a hydrogen-ion concentration of 1 x 10-5 M).

Answer: B.